OpenStudy (anonymous):

A number is increased by 80%, or 480 . What is the original number?

Parth (parthkohli):

So that means 180% of your number is 480. Can you translate that into an equation?

OpenStudy (mathstudent55):

Let the number = x x + 80%x = 480 x + 0.8x = 480 1.8x = 480 Solve the equation for x.

OpenStudy (anonymous):

4367

Parth (parthkohli):

Really?

OpenStudy (anonymous):

Yes.

Parth (parthkohli):

Hmm, I don't think so.

OpenStudy (anonymous):

267

Parth (parthkohli):

Well, \(267 \times 1.8 = 480.6\)

OpenStudy (anonymous):

MODA 400/

OpenStudy (anonymous):

400,600, OR 2400

OpenStudy (anonymous):

@openstudy

OpenStudy (anonymous):

Does the owner of this site have an account or no

hero (hero):